When answering, please refer to your answers according to the NUMBER in the box. Need to know answer but also if there was change or no change in column #9

Finding Equilibrium for I = 50, G = 50, and T = 50

(1)

Output

(Income)

Y

(2)

Net

Taxes

T

(3)

Disposable

Income

YdY-T

(4)

Consumption

Spending

C = 100+ .25Yd

(5)

Saving

Yd-C

(6)

Planned

Investment

Spending

I

(7)

Government

Purchases

G

(8)

Planned

Aggregate

Expenditure

C+I+G

(9)

Unplanned

Inventory

Change

Y-(C+I+G)

300

1

2

3

4

5

6

7

8

500

9

10

11

12

13

14

15

16

700

17

18

19

20

21

22

23

24

900

25

26

27

28

29

30

31

32

1,100

33

34

35

36

37

38

39

40

1,300

41

42

43

44

45

46

47

48

1,500

49

50

51

52

53

54

55

56
Answer #1

C = 100+ .25Yd

I = 50, G = 50, and T = 50

As it has been given above information in the table.

Hence in equilibrium

Y=C+I+G

Y=100+0.25Yd+50+50

Y=200+0.25(Y-T)

Y=200+0.25(Y-50)

Y=200+0.25Y-12.5

Y-0.25Y=187.5

0.75Y=187.5

Y=187.5/0.75

=$199.25

Hence equilibrium income level is=$199.25